A garden has a shape of an isosceles right triangle. The length of the hypotenuse is 24 feet. What is the Area of the garden?
Answer by mananth(16949)

Pythagoras theorem
leg1^2+leg2^2= hypotenuse ^2
let the legs be x each
x^2+x^2= 24^2
2x^2= 24^2
x^2= 24 *24/2
x^2= 24*12
x^2= 288
Area of right triangle = 1/2 leg^2
=1/2 *288
= 144 sq. ft